# Demonstration of a Field-Effect Three-Terminal Electronic Device with an Electron Mobility Exceeding 40 Million cm2/(Vs)

### v2

@@ Abstract
For many years, the electron mobility in the cleanest GaAs/AlGaAs heterostructures plateaued in the 30 cm2/(Vs) range. While several important discoveries were made on pristine materials (see, e.g., [6]), fabricating ultra-high electron mobility field-effect devices remained challenging. Due to various factors detrimental to 2DEG quality, field-effect-gated devices exhibited post-fabrication mobilities of no more than cm2/(Vs) [7], despite access to higher-mobility materials. Very recently, an important advance in GaAs/AlGaAs MBE growth and engineering led to 2DEGs with electron mobility reaching cm2/(Vs) [8], breaking the previous world record significantly. Nevertheless, the persistent challenge of mobility degradation during fabrication remains a critical bottleneck. This work presents a solution: a flip-chip fabrication method in which the GaAs/AlGaAs wafer undergoes no processing. Here, we demonstrate a source-drain-gate three-terminal field-effect device with electron mobility exceeding cm2/(Vs), with no measurable degradation from the pristine material.
@@
This advance is likely to impact field-effect devices broadly, enabling the fabrication of quantum wires, dot or anti-dot arrays, and electron interferometers with unprecedented ultra-high electron mobility. Moreover, our approach extends beyond GaAs/AlGaAs and could also be applied to other 2DEG materials, such as exfoliated graphene or future systems that may surpass GaAs/AlGaAs in performance. While engineering studies are ongoing, this method shows promise for low-temperature electronic devices of exceptional quality, including potentially HEMT amplifiers with superior electron mobility – of interest for the growing field of cryogenic quantum computation.

### v3

@@ Abstract
For many years, the electron mobility in the cleanest GaAs/AlGaAs heterostructures plateaued in the 30 cm2/(Vs) range. While several important discoveries were made on pristine materials (see, e.g., [6]), fabricating ultra-high electron mobility field-effect devices remained challenging. Due to various factors detrimental to 2DEG quality, and within one spectacular exception recently achieved in graphene [7], field-effect-gated devices exhibited post-fabrication mobilities of no more than cm2/(Vs) [8] despite having access to higher-mobility materials. Very recently, an important advance in GaAs/AlGaAs MBE growth and engineering led to 2DEGs with electron mobility reaching cm2/(Vs) [9], breaking the previous world record significantly. Nevertheless, the persistent challenge of mobility degradation during fabrication remains a critical bottleneck. This work presents a solution: a flip-chip fabrication method in which the GaAs/AlGaAs wafer undergoes no processing. Here, we demonstrate a source-drain-gate three-terminal field-effect device with electron mobility exceeding cm2/(Vs), with no measurable degradation from the pristine material.
@@
This advance is likely to impact field-effect devices broadly, enabling the fabrication of quantum wires, dot or anti-dot arrays, and electron interferometers with unprecedented ultra-high electron mobility and could extend beyond GaAs/AlGaAs and be applied to other 2DEG materials. In this spirit of fabricating electronic devices with unprecedented quality, we mention recent work in graphene [7] where electron mobility exceeding 50 million cm2/ were fabricated using proximity gates, albeit with electron density much lower than in our work, i.e. in the cm-2 range. Our work could also potentially proved important for future materials platforms that could surpass current state-of-the-art 2DEGs in terms of low disorder and device performance. While engineering studies are ongoing, this method shows promise for low-temperature electronic devices of exceptional quality, including potentially HEMT amplifiers with superior electron mobility – of interest for the growing field of cryogenic quantum computation.